OpenStudy (anonymous):

According to the AMERICAN PODIATRIC MEDICAL ASSOCIATION, the bones in your foot account for one-fourth of all the bones in your body.Express the number of bones in your foot in terms of the total number of bones in your body.

OpenStudy (unklerhaukus):

\[\frac{\text{ bones in foot}}{\text{bones in body}}=\frac 14\] \[\text{ bones in foot}=\frac 14 \text{bones in body}\]

OpenStudy (anonymous):

I don't get it